OpenStudy (anonymous):

The circumference of the Earth is approximately 25,000 miles at the equator. What is the distance in scientific notation?

OpenStudy (anonymous):

2.5 X 10^4

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@gorv

OpenStudy (anonymous):

@Patrickweewoos

OpenStudy (triciaal):

general form for scientific notation a * 10^ n a > 0 and a< 10 n is an integer if you move the decimal left you are dividing you you multiply to keep the same value if you move the decimal right you are multiplying so you divide to keep the same value n is the number of places you move the decimal point

OpenStudy (anonymous):

Thank yoiu

OpenStudy (triciaal):

example 34.6 = 3.46 * 10^ 1 example .00257 = 2.57 * 10^-3